Design 30 network from this address book:195.5.2.0/24


1
Expert's answer
2021-06-13T10:23:39-0400

IP Address: 195.5.2.0/24

Subnet Mask 255.255.255.0


Mask in binary 1111111.11111111.11111111.00000000


Number of possible subnets: 2N , where N equals the number of host bits borrowed.

To get 30 or more subnets we need to borrow 5 host bits from the fourth octet


25 = 32 subnets


New subnet mask in binary 11111111.11111111.11111111.11111000 or /29 in CDIR

We will use any of the 30 subnets

Number of hosts for each subnet will be

2H – 2 = 23 -2 = 6 hosts per subnet


The subnet list order

Subnet number

Subnet ID

Valid host range, subnet broadcast IP


1

195.5.2.0

195.5.2.1 - 195.5.2.6

195.5.2.7


2

195.5.2.8

195.5.2.9 - 195.5.2.14

195.5.2.15


3

195.5.2.16

195.5.2.17 - 195.5.2.22

195.5.2.23


4

195.5.2.24

195.5.2.25 - 195.5.2.30

195.5.2.31


5

195.5.2.32

195.5.2.33 - 195.5.2.38

195.5.2.39


6

195.5.2.40

195.5.2.41 – 195.5.2.46

195.5.2.47


7

195.5.2.48

195.5.2.49 - 195.5.2.54

195.5.2.55


8

195.5.2.56

195.5.2.57 - 195.5.2.62

195.5.2.63


9

195.5.2.64

195.5.2.65 - 195.5.2.70

195.5.2.71


10

195.5.2.72

195.5.2.73 - 195.5.2.78

195.5.2.79


11

195.5.2.80

195.5.2.81 - 195.5.2.86

195.5.2.87


12

195.5.2.88

195.5.2.89 - 195.5.2.94

195.5.2.95


13

195.5.2.96

195.5.2.97 - 195.5.2.102

195.5.2.103


14

195.5.2.104

195.5.2.105 - 195.5.2.110

195.5.2.111


15

195.5.2.112

195.5.2.113 - 195.5.2.118

195.5.2.119


16

195.5.2.120

195.5.2.121 - 195.5.2.126

195.5.2.127


17

195.5.2.128

195.5.2.129 - 195.5.2.134

195.5.2.135


18

195.5.2.136

195.5.2.137 - 195.5.2.142

195.5.2.143


19

195.5.2.144

195.5.2.145 - 195.5.2.150

195.5.2.151


20

195.5.2.152

195.5.2.153 - 195.5.2.158

195.5.2.159


21

195.5.2.160

195.5.2.161 - 195.5.2.166

195.5.2.167


22

195.5.2.168

195.5.2.169 - 195.5.2.174

195.5.2.175


23

195.5.2.176

195.5.2.177 - 195.5.2.182

195.5.2.183


24

195.5.2.184

195.5.2.185 - 195.5.2.190

195.5.2.191


25

195.5.2.192

195.5.2.193 - 195.5.2.198

195.5.2.199


26

195.5.2.200

195.5.2.201 - 195.5.2.206

195.5.2.207


27

195.5.2.208

195.5.2.209 - 195.5.2.214

195.5.2.215


28

195.5.2.216

195.5.2.217 - 195.5.2.222

195.5.2.223


29

195.5.2.224

195.5.2.225 - 195.5.2.230

195.5.2.231


30

195.5.2.232

195.5.2.233 - 195.5.2.238

195.5.2.239


31

195.5.2.240

195.5.2.241 - 195.5.2.246

195.5.2.247


32

195.5.2.248

195.5.2.249 - 195.5.2.254

195.5.2.255